Use the information provided to write the standard form equation of the parabola.Vertex:(3,4) Focus:(3 ¹⁵/⁴)


1
Expert's answer
2021-11-24T10:06:32-0500

If you have the equation of a parabola in vertex form "y=a(x\u2212h)^2+k," then the vertex is at "(h,k)" and the focus is "(h, k+1\/(4a))"

Vertex: "(3,4)." Then "h=3, k=4."

Focus: "(3, 15\/4)." Then "4+1\/(4a)=15\/4=>a=-1"


The equation of a parabola in vertex form is "y-4=-(x-3)^2."

Or

"y=-x^2+6x-5"